### v1

# The Quest for the Origins of Ultra-High-Energy Cosmic Rays

Abstract
Significant progress has been made over the past decades towards unveiling the sources of the most energetic particles in nature, the ultra-high-energy cosmic rays (UHECRs). Despite these advancements, the exact astrophysical sites capable of accelerating these particles to such extreme energies remain largely unknown. Moreover, the mechanisms by which they achieve these extreme energies are poorly understood. Here, I provide a concise overview of the theory underlying the acceleration and propagation of ultra-high-energy cosmic rays. I then critically discuss three recent results that could help unveil their origins: the reported excess around Centaurus A, the correlation with starburst galaxies, and the efforts to jointly model the energy spectrum, composition, and arrival directions. Finally, I discuss strategies for advancing this field, emphasising the need for refined theoretical models, the challenges in building them, and the potential for new observatories to shed light on the mysteries of UHECRs.
keywords:
ultra-high-energy cosmic rays - cosmic-ray acceleration - cosmic-ray propagation - high-energy astrophysical sources
1 Introduction
[…]
To solve these central issues, it is important to carefully consider how to interpret the data collected by UHECR observatories. This task is not trivial, as it requires highly elaborate models of how UHECRs are produced, how they escape the environments surrounding their sources, and how they propagate through the universe. This, evidently, requires sophisticated computational tools for model-building, in addition to advanced statistical methods to make the most of the available data.
In this work, I focus on the state-of-the-art methods and theoretical frameworks used to interpret UHECR data, with emphasis on the limitations of current models and the challenges that remain. I begin in section §2 by briefly presenting how UHECRs are accelerated to such extreme energies, and the key observables that are used to study them. In §3, I examine the propagation of UHECRs in the universe. Section §4 is dedicated to a theoretical discussion of some interesting recent observations.
2 UHECR sources and acceleration
The mechanisms whereby cosmic rays can obtain ultra-high energies are not fully understood. Conceptually, one could think of two ways for CRs to acquire their extreme energies, depending on whether they start with even higher energies and end up with the energies we detect – the so-called top-down models – or whether they are accelerated to the observed energies – the bottom-up models.
Bottom-up models are a more orthodox description of particle acceleration. The archetypical case is electromagnetic acceleration, wherein electromagnetic processes in the vicinity of their sources donate energy to the CRs. This includes a variety of acceleration templates such as the well-known second-order Fermi acceleration (Fermi, , 1949), diffusive shock acceleration (Bell, 1978a, ; Bell, 1978b, ; Drury, , 1983), magnetic reconnection (de Gouveia Dal Pino and Lazarian, , 2000; Giannios, , 2010), unipolar induction (Goldreich and Julian, , 1969; Lovelace, , 1976), one-shot mechanisms (Caprioli, , 2015), shear acceleration (Berezhko, , 1981; Berezhko and Krymskii, , 1981), and many others (Matthews et al., , 2020). The second type is gravitational acceleration, wherein particles are accelerated by the gravitational field of massive objects, such as black holes (Bañados et al., , 2009; Zaslavskii, , 2010, 2012). Despite its low popularity and the criticisms it has received, this type of acceleration remains a valid option until it can be fully tested (see Harada and Kimura, , 2014, for a review).
Top-down models are based on the decay of supermassive particles, such as topological defects or superheavy dark matter particles, which would produce UHECRs. Intuitively, top-down models tend to create more problems than they solve, as they require an explanation for how such highly energetic particles were created in the first place. The non-observation of UHE photons provides evidence against several such models, imposing stringent constraints on their viability (Telescope Array Collaboration, , 2019; Pierre Auger Collaboration, 2022a, ; Pierre Auger Collaboration, 2023b, ).
2.1 Some criteria for UHECR acceleration
Under the electromagnetic acceleration paradigm (and partly for the gravitational acceleration one), UHECR sources generally satisfy a few constraints (Ptitsyna and Troitsky, , 2010), listed below.
$\triangleright$ Geometrical constraints. A useful geometrical criterion commonly used to identify possible UHE accelerators was identified by Hillas, (1984). The so-called Hillas condition represents a necessary (but not sufficient) condition for particle acceleration. It states that the size of the confining region ($R_{\text{o}}$) must be larger than the Larmor radius of the particle ($R_{\text{L}}$):
$R_{\text{o}}\gtrsim R_{\text{L}}\;\;\Rightarrow\;\;R_{\text{o}}\gtrsim\dfrac{1}{Z}\left(\dfrac{B_{\text{o}}}{1\;\text{nT}}\right)^{-1}\left(\dfrac{E}{10\;\text{EeV}}\right)\;\text{kpc}\,,$ (1)
where $B_{\text{o}}$ is the magnetic field in the acceleration region, $E$ is the energy of the particle, and $Z$ is the atomic number of the CR nucleus. A schematic illustration of the Hillas criterion is shown in fig. 1. The corresponding maximum energy achievable by a given source does not follow immediately from eq. 1, as it depends on the typical velocity of the fastest particles. In this case, the geometrical condition given by eq. 1 leads to the following dynamical constraint on the maximal energy ($E_{\text{max}}$):
$E_{\text{max}}\lesssim Z\beta R_{\text{o}}B_{\text{o}}\;\;\Rightarrow\;\;E_{\text{max}}\lesssim 10Z\beta\left(\dfrac{B_{\text{o}}}{1\;\text{nT}}\right)\left(\dfrac{R_{\text{o}}}{1\;\text{kpc}}\right)\;\text{EeV}\,.$ (2)
The proportionality of $E_{\text{max}}$ to $Z$ implies a rigidity dependence $R_{\text{max}}=E_{\text{max}}/Ze$. Note that, for relativistic shocks, a Lorentz factor ($\gamma$) can increase $E_{\text{max}}$ from eq. 2. In reality, the acceleration process is much more complex than what is described by the Hillas criterion. Energy losses and magnetic field inhomogeneities can significantly affect the acceleration process. Interested readers are directed to the review by Matthews et al., (2020).
Figure 1: The Hillas diagram, illustrating different classes of possible particle accelerators. The vertical axis represents the magnetic field strength of the confining region, whereas the horizontal refers to the size of this region, in the co-moving frame of the source. Solid lines indicate the energies above which particles can no longer be trapped by the object, under the assumption of Bohm diffusion. The dashed lines indicate the energies corresponding to the knee and ankle features of the CR spectrum. The information used in this plot is based on reviews by Alves Batista et al., 2019a and Anchordoqui, (2019).
$\triangleright$ Power constraints. The energy budget available for acceleration at any given instant must be sufficient to be used for acceleration. Assuming that this energy is supplied by the magnetic field, one arrives at the Hillas-Lovelace limit (Lovelace, , 1976; Hillas, , 1984):
$P_{B,\text{min}}\gtrsim u_{B}R_{\text{o}}^{2}v\;\;\Rightarrow\;\;P_{B,\text{min}}\gtrsim\dfrac{10^{37}\eta^{2}}{\beta Z^{2}}u_{B}\left(\dfrac{E_{\text{max}}}{10\;\text{EeV}}\right)^{2}\;\text{J}\,\text{s}^{-1}\;,$ (3)
where $\eta$ refers to an efficiency factor, and $u_{B}$ the magnetic energy density. Eq. 3 is oversimplified here and holds for the non-relativistic case. In reality, the power $P_{B,\text{min}}$ depends on details of the acceleration process and on the diffusion properties of the medium (for more detailed considerations, see, e.g., Waxman, , 1995; Matthews et al., , 2020; Matthews and Taylor, , 2023).
$\triangleright$ Energy gain vs. losses. The energy gained by particles during acceleration must necessarily exceed their total energy losses, including radiative losses and interactions with the surrounding medium. Information about the source environment, such as the density of matter and radiation in the medium as well as the magnetic field strength, can be effectively used to rule out potential source candidates.
Ptitsyna and Troitsky, (2010) considered two additional conditions. The first is related to the total emissivity of a given source population, whose number density must be sufficient to account for the observed flux of UHECRs. While valid, this condition is weaker than the others, as it is a posteriori, depending strongly on propagation details which are not well known (see §3). For example, it presumes some level of knowledge about the source population, including temporal characteristics (steady emission vs. bursts), among other factors. The second additional condition involves multi-messenger signatures and is also a posteriori. It requires that emissivity constraints be consistent with observational limits on the diffuse fluxes of lower-energy CRs, electromagnetic radiation, and neutrinos. This condition is stronger than the aforementioned density-emissivity limits, as it incorporates information from multiple independent probes (messengers).
3 UHECR propagation
The propagation of UHECRs through intergalactic space is a complex process that depends on the properties of the particles themselves, the magnetic fields they traverse, and the background radiation fields with which they interact (see §3.1 and §3.2). Accurately modeling UHECRs propagation requires accounting for these factors, which are typically combined within computational tools designed to simulate their transport (see §3.3).
3.1 Interactions with matter and radiation
The interaction of UHECRs with matter is generally considered negligible due to the low density of gas in the intergalactic medium (IGM). However, it can be important in other environments (Owen et al., , 2018), playing a particularly special role in gamma-ray and neutrino emission by galaxy clusters (Fang and Olinto, , 2016; Hussain et al., , 2021, 2023).
One of the main interactions of UHECRs with photon fields is photomeson production – or photopion production, since pion production is by far the dominant channel –, which is the interaction of CR nuclei with pervasive radiation fields predominantly of the cosmic microwave background (CMB), extragalactic background light (EBL), and cosmic radio background (CRB). The well-known Greisen-Zatsepin-Kuz’min (GZK) cut-off (Greisen, , 1966; Zatsepin and Kuz’min, , 1966) is exactly the process of UHE protons interacting with CMB photons leading to a strong suppression for energies above $E_{\text{GZK}}\approx 5\times 10^{19}\;\text{eV}$. UHECR interactions with photons can incur energy losses due to the production of electron-positron pairs, a process known as Bethe-Heitler pair production (Bethe and Heitler, , 1934). Nuclei heavier than hydrogen can also undergo photodisintegration. Together, all these processes can significantly affect the spectrum of UHECRs. The energy loss lengths associated with these interactions, along with adiabatic energy losses due to the expansion of the universe, are illustrated in fig. 2.
Figure 2: Energy loss length as a function of the energy, at $z=0$. Different types of nuclei are represented by distinct colours. Thin faint lines with varying styles refer to specific photonuclear process. The thick solid lines show the total contribution from all processes. Results are based on the EBL model by Saldana-Lopez et al., (2020), derived using the CRPropa code (Alves Batista et al., , 2022).
Propagation uncertainties in UHECRs were investigated by Alves Batista et al., (2015), who examined the effects of EBL models and photodisintegration cross sections. Their study demonstrated that these factors can significantly influence the observed spectrum and composition. Soriano et al., (2018) conducted a detailed analysis of ${}^{4}\text{He}$ photodisintegration, revealing that its cross section had been overestimated compared to current measurements. Notably, photodisintegration cross sections also play a critical role in understanding the escape of CRs from their sources (Boncioli et al., , 2017). Additional challenges arise when considering photomeson production for heavier nuclei, as shown by Morejon et al., (2019).
3.2 Magnetic fields
The presence of extragalactic magnetic fields can significantly influence the propagation of UHECRs. Naturally, the EGMF between sources and Earth determines by how much CRs are deflected and how much longer they must travel before reaching Earth. Since most of the universe’s volume consists of “empty” cosmic voids, it is reasonable to expect that the fields in these regions dominate the propagation of UHECRs, although this remains to be proven. This is captured by the notion of volume filling factor, illustrated in fig. 3. One could, of course, argue that outflows from large-scale structures like galaxy clusters can inject magnetised material into the IGM, but this effect should not significantly alter the magnetisation of cosmic voids over the age of the universe (e.g., Furlanetto and Loeb, , 2001; Barai et al., , 2008; Beck et al., , 2013; Arámburo-García et al., , 2021). Therefore, unless sources are located very close to Earth such that they do not traverse cosmic voids, UHECRs would likely cross these regions at least once.
Figure 3: Cumulative volume filling factors for the EGMF models by Hackstein et al., (2018) (green lines) and the upper-limit models of Alves Batista et al., (2017) (orange lines) at $z=0$. Each line style represents a different initial condition for the magnetic field, as described in the corresponding references.
To a first approximation, neglecting interactions and energy losses, the deflection of UHECRs depend on the distance from Earth to the source relative to the coherence length of the field ($L_{B}$), yielding:
$\delta\simeq\begin{cases}5.3^{\circ}Z\left(\dfrac{D}{1\;\text{Mpc}}\right)\left(\dfrac{B_{\perp}}{100\;\text{fT}}\right)\left(\dfrac{E}{10\;\text{EeV}}\right)^{-1}&\text{if}\;D\ll L_{B}\,,\\ 1.7^{\circ}Z\left(\dfrac{D}{1\;\text{Mpc}}\right)^{\frac{1}{2}}\left(\dfrac{B_{\perp}}{100\;\text{fT}}\right)\left(\dfrac{L_{B}}{100\;\text{kpc}}\right)^{\frac{1}{2}}\left(\dfrac{E}{10\;\text{EeV}}\right)^{-1}&\text{if}\;D\gg L_{B}\,.\\ \end{cases}$ (4)
[…]
3.3 Propagation codes
Propagation codes such as CRPropa (Armengaud et al., , 2007; Kampert et al., , 2013; Alves Batista et al., , 2016, 2022), SimProp (Aloisio et al., , 2017), PriNCe (Heinze et al., , 2019), and TransportCR (Kalashev and Kido, , 2015) are widely used to model the propagation of UHECRs in the universe. CRPropa and SimProp rely on Monte Carlo methods, whereas PriNCe and TransportCR employ transport equations. CRPropa also offers a three-dimensional mode, including magnetic-field effects.
None of these codes can capture the full complexity of UHECR propagation within a feasible time frame, leaving much of the parameter space unexplored. Unfortunately, this will likely not change in the near future, particularly given the stochastic nature of some aspects of propagation, which makes any attempts to invert this problem seemingly unfeasible. Unless we are fortunate enough that all relevant UHECR parameters lie within the small region of parameter space accessible to current methods, new techniques will be essential to advance the field. For more in-depth discussion, see the white paper of the European Consortium for Astroparticle Theory (EuCAPT) (Alves Batista et al., , 2021).
3.4 Cosmogenic particles
An important consequence of the extragalactic propagation of UHECRs is the production of cosmogenic particles, namely neutrinos and photons. In particular, the decay of neutral ($\pi^{0}\rightarrow\gamma+\gamma$) and charged ($\pi^{+}\rightarrow\mu^{+}+\nu_{\mu}\rightarrow e^{+}+\nu_{e}+\bar{\nu}_{\mu}+\nu_{\mu}$) pions produced via photopion production, together with beta decays ($n\rightarrow p+e^{-}+\bar{\nu}_{e}$ ) generated during photodisintegration, abundantly produce neutrinos and fuel intergalactic electromagnetic cascades initiated by photons and electrons.
An interesting aspect of cosmogenic neutrinos is that they can be used to constrain the fraction of protons in the UHECR flux, for a given source emissivity (van Vliet et al., , 2019; Muzio et al., , 2023). However, current estimates of the flux of cosmogenic neutrinos suggest a rather bleak scenario for their detection (Alves Batista et al., 2019b, ; Heinze et al., , 2019). But given the limitations of the phenomenological models employed (see §4.3), this pessimism might not be warranted. Regarding the flux of cosmogenic photons, its contribution to the diffuse gamma-ray background (DGRB) at energies of $\sim 1\;\text{TeV}$ may strongly constraint some UHECR models (Alves Batista et al., 2019b, ; van Vliet et al., , 2017).
4 Interpreting the observations
The primary astrophysical observables derived from UHECR data are the energy spectrum, mass composition, and arrival directions. They are essential for unraveling the origin and propagation of UHECRs, as well as for differentiating between competing source models. In the following subsections, I highlight a few relevant findings from the last decade to illustrate the state of the field, and critically examine what we have learned so far.
4.1 Centaurus A
One of the most compelling results from the last decade is the correlation of UHECRs and the nearest active galactic nucleus (AGN), Centaurus A (Cen A), located at approximately 3.8 Mpc from Earth. This radio-loud AGN has been long deemed a strong candidate to accelerate CRs to UHE (Romero et al., , 1996). The central galaxy itself hosts a supermassive black hole of approximately $55\times 10^{6}M_{\odot}$, and contains a pair of radio jets, likely powered by accretion onto the central black hole, extending over several kiloparsecs (Hardcastle et al., , 2003; Müller et al., , 2011, 2014).
Several suggestions for how UHECRs could be produced in Cen A have been put forward. The way in which particles are accelerated and the region where this happens vary hugely depending on the model, possibly taking place in the jet itself, the radio lobes, its inner core, and even in jet-lobe interactions (Honda, , 2009; Kachelrieß et al., , 2009; Fraija, , 2014; McKinley et al., , 2018; Rodríguez-Ramírez et al., , 2019). Cen A is also a starburst galaxy (SBG), displaying signs of intense star formation activity, possibly due to its history of galaxy mergers (Crockett et al., , 2012).
The Pierre Auger Collaboration, 2022b detected an anisotropic distribution of UHECRs at energies $E>38\;\text{EeV}$. A significant fraction of these CRs are clustered around the direction of Cen A, within a $27^{\circ}$ window. The significance of this excess is $\approx 3.9\sigma$, depending on details of the analysis (Pierre Auger Collaboration, , 2018).
If the Cen A signal continues to grow over time, by 2026 it could reach a $5\sigma$ significance. Suppose this is the case: could we claim that Cen A is a source of UHECRs? The answer is unequivocally no. From a theoretical viewpoint, the necessary conditions to assert that Cen A (or any other object) is an UHECR source are:
  • The energy budget available should be sufficient to generate the observed flux of UHECRs within a given time window.
  • If acceleration is electromagnetic, the size of the confining magnetised region must enable the escape of CRs with the energy of interest.
  • The environment where acceleration occurs must sufficiently transparent to UHECRs, allowing their escape. Testing this necessarily requires multi-messenger and multi-wavelength observations.
  • The source must be unambiguously localised, meaning it has to be convincingly distinguisheable from other sources located in the vicinity or along the same line of sight.
[…]
    EGMF between the Milky Way and the object must be either reasonably well-known, to enable a precise identification of the object’s coordinates in the sky, or they ought to be sufficiently small to ensure a high-confidence identification of the emitting region.
  • The GMF must be either sufficiently well-known to be accurately modelled, or weak enough that its impact on CRs is negligible.
These conditions are all necessary for identifying any UHECR source, Cen A included, but they are not sufficient.
4.2 The starburst correlation
Starburst galaxies are considered potential sources of UHECRs due to their intense star formation activity, which creates an environment conducive to particle acceleration. Their magnetic fields, several orders of magnitude stronger than that of the Milky Way, enable longer confinement times, thereby allowing acceleration to higher energies (Völk et al., , 1996; Bykov, , 2014). However, the highly turbulent and dense interstellar medium (ISM) in SBG could, in principle, absorb a significant fraction of the CRs produced within them. A way to test this hypothesis is by searching for high-energy neutrinos from SBG (e.g., Loeb and Waxman, , 2006; Lunardini et al., , 2019; Ambrosone et al., , 2021). Gamma rays, which can be produced by the same mechanisms as neutrinos, might also serve as a counterpart (H.E.S.S. Collaboration, , 2009; Bykov, , 2014; Ohm, , 2016; Linden, , 2017), but the ISM could also be opaque to their propagation, making the detection of this counterpart more challenging.
The high supernova rate in SBG likely contributes to a significantly higher flux of high-energy CRs compared to normal galaxies. Additionally, winds and outflows driven by star formation are potential accelerators (e.g, Anchordoqui, , 2018), especially when pre-accelerated particles from supernova (SN) shocks are considered.
Currently a $4\sigma$ evidence of correlation between UHECRs and SBG has been reported by Auger (Pierre Auger Collaboration, , 2018; Pierre Auger Collaboration, 2022b, ). If this signal keeps increasing over time, can we unambiguously claim that SBG are sources of UHECRs? The answer is: it depends what is meant by sources. Given that this association is true and that there are no sources along the same line of sight, it is possible to say that UHECRs come from SBG, but not that the SBG themselves are sources. The necessary conditions for them to be sources are the same as for Cen A. But there is an additional issue that deserves attention: given the high star formation rate in these galaxies, it is well possible that some objects such as young pulsars (Blasi et al., , 2000; Fang et al., , 2012) or magnetars (Arons, , 2003) are the true sources.
Anchordoqui, (2018) proposed that correlation with all matter in the local universe would favour the interpretation that the sources are embedded in SBG. He also argues that this is not viable due to energy-budget considerations, and puts forward a model based on SBG winds. Since Auger’s results suggest that at least $\sim 10\%$ of the total UHECR flux comes from SBG, it is rather puzzling that Telescope Array (TA) has not observed a significant correlation (Telescope Array Collaboration, 2018b, ).
What is more: even if we can unambiguously confirm that UHECRs come from SBG, we are still not able to identify the acceleration mechanism. As described in §2, acceleration mechanisms are simply templates and can be applied to a variety of types of objects which could also contribute to the observed flux. Then the question becomes: what can we learn from the UHECRSBG correlation? On its own, at present, we have not learnt anything fundamentally new. However, this signal is a first step toward unveiling the sources of UHECRs. With higher statistics and better knowledge of propagation ingredients such magnetic fields and background photon fields, it is possible to add information from other messengers, in particular neutrinos and gamma rays, to build a self-consistent model that describes UHECR emission by SBG. Note that multi-messenger and multi-wavelength information are necessary conditions to meaningfully interpret these observations but they are still not sufficient.
In summary, to understand the apparent UHECRSBG correlation, one needs to be able to disentangle between the following scenarios:
  1. (i)
    astrophysical objects in SBG that accelerate UHECRs;
  2. (ii)
    large-scale phenomena inherent to SBG themselves, such as super-winds, are the accelerators;
  3. (iii)
    a combination of the two, with pre-acceleration in objects and some appreciable acceleration by large-scale phenomena afterwards.
4.3 Combined spectrum-composition fits
A notable surge in the number of attempts to perform combined spectrum-composition fits has been observed in the last decade (Hooper and Taylor, , 2010; Taylor et al., , 2014, 2015; Pierre Auger Collaboration, 2017a, ; Alves Batista et al., 2019b, ; Heinze et al., , 2019; Pierre Auger Collaboration, 2023a, ; Pierre Auger Collaboration, 2024a, ; Pierre Auger Collaboration, 2024b, ). These fits are phenomenological, starting with an injection function ($Q_{\text{src}}$), defined as the number of particles of a given nuclear species emitted by a source, per energy, at a given time interval
$Q_{\text{src}}\left(E,t\right)\equiv\dfrac{\text{d}^{2}N}{\text{d}E\,\text{d}t}=\sum\limits_{\varkappa\in\kappa}\mathcal{S}_{\varkappa}(E,t;E_{\text{cut}}^{(\varkappa)}(t))\;,$ (5)
where $\kappa=\{\left(A,Z\right)\;|\;\left(A>0,\,Z\geq 0\right)\}$ is the set of all CR species emitted, with atomic number $Z$ and atomic mass $A$. Here $E$ is the energy of the emitted CR, $\mathcal{S}_{\varkappa}$ is a function that describes the spectrum, with a characteristic composition-dependent maximal energy ($E_{\text{cut}}^{(\varkappa)}$), and $t$ is a time variable, which is related to the redshift.
A common assumption is that UHECR acceleration is electromagnetic and that the cut-off energy scales linearly with the particle charge ($Ze$), as argued by Peters, (1961). Even though this assumption is reasonable and serves as a useful starting point, astrophysical environments can contain matter and radiation fields that affect $\mathcal{S}_{\varkappa}$ (Unger et al., , 2015).
$\mathcal{S}_{\varkappa}$ is commonly decomposed into a power law of the form $E^{-\alpha_{\varkappa}}$, wherein $\alpha_{\varkappa}$ is the spectral index of the source, and a cut-off function $\mathcal{G}(E,t;E_{\text{cut}^{(\varkappa)}}(t))$, which is often taken as an exponential cut-off:
$S_{\varkappa}(E,t;E_{\text{cut}}^{(\varkappa)}(t))=E^{-\alpha_{\varkappa}}\,\mathcal{G}(E,t;E_{\text{cut}}^{(\varkappa)}(t))\;.$ (6)
Recently, Ehlert et al., (2023) adopted a phenomenological parametrisation for the cut-off, $E_{\text{cut}}^{(\varkappa)}=E_{\text{cut}}\left(A,Z\right)=E_{0}Z^{a}A^{b}$, wherein $E_{0}$ is a reference energy. In this case, Peters’ cycle (Peters, , 1961) is a special case of this parameterisation, with $a=1$ and $b=0$, and models with in-source photodisintegration (Unger et al., , 2015) have $a=0$ and $b=1$. They have shown that the $\mathcal{G}$ plays an important role in the outcome of these fits.
$\triangleright$ Source distribution. A common assumption in combined fits is that the UHECR sources are homogeneously distributed in space. This is a reasonable assumption for UHECRs of energies $E\lesssim 10\;\text{EeV}$, whose typical energy loss lengths are comparable to or larger than the distance to the sources, such that most sources are contributing to the observed flux. However, at higher energies ($E\gtrsim 40\;\text{EeV}$), the sources are mostly local (see fig. 2), lying within the interaction horizon ($\lesssim 100\;\text{Mpc}$), which means that they are not homogeneously distributed. More recently, some progress has been made in this direction, considering specific source populations such as AGN and minding the potential importance of possible nearby sources like Cen A to the observed flux (Pierre Auger Collaboration, 2024a, ). Moreover, new analyses looking for correlations with other messengers have also been informing us on the distribution of sources (Partenheimer et al., , 2024).
$\triangleright$ Source luminosity. Frequently the sources are assumed to have the same luminosity, except for a few works that fix it with respect to intrinsic source properties (for instance, Eichmann et al., , 2018; Biehl et al., , 2018; Boncioli et al., , 2019; Eichmann et al., , 2022). This assumption is helpful to enable the interpretation of observations, but it is far from realistic. In reality, some sources are most likely brighter than others, depending on their size, age, and on details of the acceleration process.
$\triangleright$ Composition. The composition at the source is essentially the metallicity in the accelerating region, which evolves with cosmic time. Still, most works assume a time-independent composition. Although, this makes the interpretation of the data much more tractable, it is not realistic. Furthermore, the composition at the source is uncertain, and even if mean free paths are similar for a few nuclear species, the by-products of their interactions can be very different.
$\triangleright$ Extragalactic magnetic fields. Most works, for simplicity, ignore the effects of EGMFs. This could be justified, but at present we lack information to support it (see §3.2). Wittkowski for the Pierre Auger Collaboration, (2017) performed a combined spectrum-composition fit considering a “realistic” distribution of EGMFs. They found that the inclusion of EGMFs could lead to a significant change in the results of the fits. For instance, in one of their scenarios, they inferred a spectral index of $\alpha\approx 1.61$ and a maximal rigidity of $\log(R_{\text{cut}}/\text{eV})\approx 18.88$, compared to $\alpha\approx 0.61$ and $\log(R_{\text{cut}}/\text{eV})\approx 18.48$, in the case where EGMFs were ignored. More recently, the Pierre Auger Collaboration, 2024b performed a similar fit considering the effects of turbulent EGMFs statistically (with one-dimensional simulations), which would imply a magnetic horizon effect on the UHECR spectrum. Even though this model is completely unrealistic, considering present-day knowledge of EGMFs (for reviews, see e.g., Alves Batista and Saveliev, , 2021; Vachaspati, , 2021), it is a step forward compared to most studies which assume an unmagnetised universe. And most importantly, it does confirm the aforementioned trend that the best-fit spectral indices tend to increase (for an $E^{-\alpha}$ spectrum). This has important implications for understanding the transition between Galactic and extragalactic CRs (Aloisio et al., , 2012; Parizot, , 2014), as magnetic horizon effects determine where the transition occurs (see, e.g., Alves Batista and Sigl, , 2014, for a discussion).
$\triangleright$ Temporal emission profile. Phenomenological fits of the UHECR observations rarely account for the CR emission time. In fact, the implicit (and rather strong) assumption is that the sources are continuously emitting CRs, which is evidently incorrect. Many source candidates are transient or have some large but finite duty cycle. Moreover, even if the sources are steady emitters, their finite lifetime could have a significant impact on phenomenological observables. Eichmann and Kachelrieß, (2023) have shown that this could lead to a hardening of the spectrum measured from Earth, under the assumption that radio galaxies are the sources of UHECRs. Nevertheless, similar outcomes could be expected for other source populations.
Recently, the Pierre Auger Collaboration, 2024a extended the combined spectrum-composition fit by including an additional observable: anisotropies. Nevertheless, the model does not properly account for the effects of EGMFs, modelling it as a directional smear around the sources. UHECR trajectories are elongated due to magnetic deflections (see eq. 4), which raises interaction rates during propagation and can substantially affect the observed spectrum, composition, and anisotropies. Previously, Eichmann et al., (2018) had already performed a combined fit of the three observables (spectrum, composition, and arrival directions), focusing on radio galaxies. Such multi-observable fits are essential and should become the norm, given their constraining power. Yet, they inevitably open a “Pandora’s box” of modelling, drastically increasing the number of free parameters and hence complexity of the model.
5 Conclusions and Outlook
We find ourselves at a crossroads of UHECR astrophysics. Existing facilities such as the Pierre Auger Observatory and the Telescope Array are expected to continue operating for another decade, but it is unclear what will happen next (Coleman et al., , 2023). Plans are underway for new dedicated UHECR observatories like the Global Cosmic-ray Observatory (GCOS(Hörandel for the GCOS Collaboration, , 2022; Alves Batista for the GCOS Collaboration, , 2023) and POEMMA (POEMMA Collaboration, , 2021), in addition to observatories primarily designed for other messengers but which can also deliver information about UHECRs, such as the Giant Radio Array for Neutrino Detection (GRAND Collaboration, , 2020).
There is a clear need for a new generation of observatories that can provide precise measurements of the energy spectrum, mass composition, and arrival directions of UHECRs, increasing exposure and reducing uncertainties. The ability to identify mass composition on an event-by-event basis, in particular, could have a great impact on model-building (see Pierre Auger Collaboration, 2024c, , for recent advances in this direction).
Even if these new observatories are built, interpreting UHECR observations is far from trivial. Major theoretical challenges persist, and the assumptions made in interpreting the data are often unjustified and at times manifestly unrealistic. Understandably, this could be useful to make a problem tractable for a first attempt at tackling it, but it should not become the standard approach. Instead, models that progressively incorporate additional layers of complexity, besides being more robust, are more likely to provide a more accurate picture of the UHE universe.
Here I selected examples that illustrate why current models must be improved to provide a meaningful interpretation of the data, and how this can be done. The phenomenological fits described in §4.3, for example, can be promptly enhanced with existing tools, although the vastness of the parameter space of uncertainties still prevents an exhaustive scan. Still, some progress can be made without awaiting the next generation of observatories. The observations discussed in §4.1 and §4.2, on the other hand, do require next-generation instruments to enable actual discoveries.
Finally, all these efforts must be pursued within a multi-messenger framework, acknowledging the fact that no single messenger, be it CRs, photons, or neutrinos, can provide the complete picture. A synergistic approach that integrates diverse observational channels is indispensable to make progress and ultimately unveil the origins of UHECRs.
Acknowledgements
I acknowledge support from the Agence Nationale de la Recherche (ANR), project ANR-23-CPJ1-0103-01. I also thank the organisers and participants of ECRS2024, who contributed to the discussions that inspired this work.

### v2

@@
 The propagation of UHECRs through intergalactic space is a complex process that depends on the properties of the particles themselves, the magnetic fields they traverse, and the background radiation fields with which they interact (see §3.1 and §3.2).
-Accurately modeling UHECRs propagation requires accounting for these factors, which are typically combined within computational tools designed to simulate their transport (see §3.3).
+Accurately modelling UHECR propagation requires accounting for these factors, which are typically combined within computational tools designed to simulate their transport (see §3.3).
 3.1 Interactions with matter and radiation
 The interaction of UHECRs with matter is generally considered negligible due to the low density of gas in the intergalactic medium (IGM).
-However, it can be important in other environments (Owen et al., , 2018), playing a particularly special role in gamma-ray and neutrino emission by galaxy clusters (Fang and Olinto, , 2016; Hussain et al., , 2021, 2023).
+However, it can be important in other environments (Owen et al., , 2018), playing a particularly special role in gamma-ray and neutrino emission by galaxy clusters (Fang and Olinto, , 2016; Hussain et al., , 2021; Condorelli et al., , 2023; Hussain et al., , 2023).
 One of the main interactions of UHECRs with photon fields is photomeson production – or photopion production, since pion production is by far the dominant channel –, which is the interaction of CR nuclei with pervasive radiation fields predominantly of the cosmic microwave background (CMB), extragalactic background light (EBL), and cosmic radio background (CRB).
@@
 The presence of extragalactic magnetic fields can significantly influence the propagation of UHECRs.
-Naturally, the EGMF between sources and Earth determines by how much CRs are deflected and how much longer they must travel before reaching Earth.
+Naturally, the EGMF between the sources and the Earth determines by how much CRs are deflected and how much longer they must travel before reaching Earth.
 Since most of the universe’s volume consists of “empty” cosmic voids, it is reasonable to expect that the fields in these regions dominate the propagation of UHECRs, although this remains to be proven.
@@
 Therefore, unless sources are located very close to Earth such that they do not traverse cosmic voids, UHECRs would likely cross these regions at least once.
+In addition to voids, other regions of the large-scale structure of the universe also influence the propagation of UHECRs, affecting their trajectories.
+For instance, these structures can act as scattering centres (Kotera & Lemoine, , 2008), potentially causing significant deflections.
+They can also trap CRs particularly when sources are embedded within them (Alves Batista et al., , 2017; Hackstein et al., , 2018).
 Figure 3: Cumulative volume filling factors for the EGMF models by Hackstein et al., (2018) (green lines) and the upper-limit models of Alves Batista et al., (2017) (orange lines) at $z=0$.
@@
   • The energy budget available should be sufficient to generate the observed flux of UHECRs within a given time window.
-  • If acceleration is electromagnetic, the size of the confining magnetised region must enable the escape of CRs with the energy of interest.
+  • If acceleration is electromagnetic, the size of the confining magnetised region must enable the escape of CRs with the energies of interest.
   • The environment where acceleration occurs must sufficiently transparent to UHECRs, allowing their escape. Testing this necessarily requires multi-messenger and multi-wavelength observations.
@@
 However, this signal is a first step toward unveiling the sources of UHECRs.
-With higher statistics and better knowledge of propagation ingredients such magnetic fields and background photon fields, it is possible to add information from other messengers, in particular neutrinos and gamma rays, to build a self-consistent model that describes UHECR emission by SBG.
+With higher statistics and better knowledge of propagation ingredients such as magnetic fields and background photon fields, it is possible to add information from other messengers, in particular neutrinos and gamma rays, to build a self-consistent model that describes UHECR emission by SBG.
 Note that multi-messenger and multi-wavelength information are necessary conditions to meaningfully interpret these observations but they are still not sufficient.
@@
   2. (ii)
-    large-scale phenomena inherent to SBG themselves, such as super-winds, are the accelerators;
+    large-scale phenomena inherent to SBG themselves, such as super-winds, acting as accelerators;
   3. (iii)
@@
 where $\kappa=\{\left(A,Z\right)\;|\;\left(A>0,\,Z\geq 0\right)\}$ is the set of all CR species emitted, with atomic number $Z$ and atomic mass $A$.
-Here $E$ is the energy of the emitted CR, $\mathcal{S}_{\varkappa}$ is a function that describes the spectrum, with a characteristic composition-dependent maximal energy ($E_{\text{cut}}^{(\varkappa)}$), and $t$ is a time variable, which is related to the redshift.
+Here $E$ is the energy of the emitted CR, $\mathcal{S}_{\varkappa}$ is a function that describes the spectrum, with a characteristic composition-dependent maximal energy ($E_{\text{cut}}^{(\varkappa)}$), and $t$ is a time variable, which may be related to the redshift ($t=t(z)$) or to the intrinsic variability of the source.
 A common assumption is that UHECR acceleration is electromagnetic and that the cut-off energy scales linearly with the particle charge ($Ze$), as argued by Peters, (1961).
@@
 $S_{\varkappa}(E,t;E_{\text{cut}}^{(\varkappa)}(t))=E^{-\alpha_{\varkappa}}\,\mathcal{G}(E,t;E_{\text{cut}}^{(\varkappa)}(t))\;.$ (6)
-Recently, Ehlert et al., (2023) adopted a phenomenological parametrisation for the cut-off, $E_{\text{cut}}^{(\varkappa)}=E_{\text{cut}}\left(A,Z\right)=E_{0}Z^{a}A^{b}$, wherein $E_{0}$ is a reference energy.
-In this case, Peters’ cycle (Peters, , 1961) is a special case of this parameterisation, with $a=1$ and $b=0$, and models with in-source photodisintegration (Unger et al., , 2015) have $a=0$ and $b=1$.
+Recently, Muzio et al., (2024) adopted a phenomenological parametrisation for the cut-off, $E_{\text{cut}}^{(\varkappa)}=E_{\text{cut}}\left(A,Z\right)=E_{0}Z^{a}A^{b}$, wherein $E_{0}$ is a reference energy.
+In this case, the Peters’ cycle (Peters, , 1961) is a special case with $a=1$ and $b=0$, and models with in-source photodisintegration (Unger et al., , 2015) have $a=0$ and $b=1$.
 They have shown that the $\mathcal{G}$ plays an important role in the outcome of these fits.
+Moreover, variations of $\mathcal{G}$ across a source population can also play a role (Ehlert et al., , 2023).
 $\triangleright$ Source distribution.
@@
 More recently, some progress has been made in this direction, considering specific source populations such as AGN and minding the potential importance of possible nearby sources like Cen A to the observed flux (Pierre Auger Collaboration, 2024a, ).
-Moreover, new analyses looking for correlations with other messengers have also been informing us on the distribution of sources (Partenheimer et al., , 2024).
+Moreover, new analyses looking for correlations with other messengers have also been informing on the distribution of sources (Partenheimer et al., , 2024).
 $\triangleright$ Source luminosity.
@@
 Most works, for simplicity, ignore the effects of EGMFs.
-This could be justified, but at present we lack information to support it (see §3.2).
+This could be justified, but at present information is lacking to support it (see §3.2).
 Wittkowski for the Pierre Auger Collaboration, (2017) performed a combined spectrum-composition fit considering a “realistic” distribution of EGMFs.
@@
 Nevertheless, similar outcomes could be expected for other source populations.
-Recently, the Pierre Auger Collaboration, 2024a extended the combined spectrum-composition fit by including an additional observable: anisotropies.
-Nevertheless, the model does not properly account for the effects of EGMFs, modelling it as a directional smear around the sources.
+Recently, the Pierre Auger Collaboration, 2024a extended the combined spectrum-composition fit by adding another observable: anisotropies.
+Nevertheless, the model does not properly account for the effects of EGMFs, simplifying it to a simple smear around the sources.
 UHECR trajectories are elongated due to magnetic deflections (see eq.
@@
 Existing facilities such as the Pierre Auger Observatory and the Telescope Array are expected to continue operating for another decade, but it is unclear what will happen next (Coleman et al., , 2023).
-Plans are underway for new dedicated UHECR observatories like the Global Cosmic-ray Observatory (GCOS(Hörandel for the GCOS Collaboration, , 2022; Alves Batista for the GCOS Collaboration, , 2023) and POEMMA (POEMMA Collaboration, , 2021), in addition to observatories primarily designed for other messengers but which can also deliver information about UHECRs, such as the Giant Radio Array for Neutrino Detection (GRAND Collaboration, , 2020).
+Plans are underway for new dedicated UHECR observatories like the Global Cosmic-ray Observatory (GCOS(Hörandel for the GCOS Collaboration, , 2022; Alves Batista for the GCOS Collaborators, , 2023; Ahlers et al., , 2025) and POEMMA (POEMMA Collaboration, , 2021), in addition to observatories primarily designed for other messengers but which can also deliver information about UHECRs, such as the Giant Radio Array for Neutrino Detection (GRAND Collaboration, , 2020).
 There is a clear need for a new generation of observatories that can provide precise measurements of the energy spectrum, mass composition, and arrival directions of UHECRs, increasing exposure and reducing uncertainties.
@@
 Major theoretical challenges persist, and the assumptions made in interpreting the data are often unjustified and at times manifestly unrealistic.
-Understandably, this could be useful to make a problem tractable for a first attempt at tackling it, but it should not become the standard approach.
-Instead, models that progressively incorporate additional layers of complexity, besides being more robust, are more likely to provide a more accurate picture of the UHE universe.
+Understandably, this can be useful for making a problem more tractable in an initial attempt to tackle it, but it should not become the default approach.
+Instead, models that progressively incorporate additional layers of complexity are not only more robust, but also more likely to provide an accurate representation of the UHE universe.
 Here I selected examples that illustrate why current models must be improved to provide a meaningful interpretation of the data, and how this can be done.
-The phenomenological fits described in §4.3, for example, can be promptly enhanced with existing tools, although the vastness of the parameter space of uncertainties still prevents an exhaustive scan.
-Still, some progress can be made without awaiting the next generation of observatories.
+The phenomenological fits described in §4.3, for example, can be promptly enhanced with existing tools.
+Howerver, the vast parameter space of uncertainties still prevents a comprehensive exploration.
+Even so, some progress can be made without waiting for the next generation of observatories.
 The observations discussed in §4.1 and §4.2, on the other hand, do require next-generation instruments to enable actual discoveries.
@@
 I acknowledge support from the Agence Nationale de la Recherche (ANR), project ANR-23-CPJ1-0103-01.
+I thank Michael Unger for useful comments on this manuscripts.
 I also thank the organisers and participants of ECRS2024, who contributed to the discussions that inspired this work.